OpenStudy (anonymous):

− 1/6 [3x + 4(y − 10y)]

OpenStudy (stacey):

You do not state the directions for this, but if you wish to simplify, start by distributing the 4 through the parenthesis.

jhonyy9 (jhonyy9):

-1/6 (3x+4(-9y))= -1/6 (3x-36y)= (-3x)/6 +(36y)/6 = (-x/2) +6y